What is the nth term of the geometric sequence that has a common ratio of 6 and 24 as its third term?
Answer by khai(18)

r = 6
3th = 24
nth term = a(r)^(n-1) {a is the first term}
you have to find the first term first.....
24 = a(6)^(3-1)
24 = 36a
a= 2/3 {so this is the first term, now you can find nth term}
nth term = (2/3)(6)^(n-1)